# Robust Federated Learning Over the Air: Combating Heavy-Tailed Noise with Median Anchored Clipping

### v1

@@ I Introduction
Federated learning (FL)[1, 2, 3, 4] is an emerging paradigm for collaborative data processing, enabling clients to benefit from high-quality model services while safeguarding the confidentiality of their private data. Nevertheless, significant challenges persist during the execution process. The frequent transmission of model information between clients and server consumes substantial network bandwidth, while the aggregation of a large number of parameters requires extensive computing resources. Moreover, although FL avoids directly aggregating user data, the exchange of model parameters could still pose risks to user privacy, particularly through inference attacks[5].
@@
A viable solution to this problem is by integrating over-the-air (OTA) computation[6, 7, 8, 9] into the FL system, leveraging the superposition property of a multiple-access channel to automatically aggregate the clients’ gradient, significantly enhancing channel utilization while concurrently reducing computational overhead[10]. Furthermore, as the server receives aggregated gradients instead of individual ones from clients [11], the vulnerability to inference attacks is significantly reduced.
@@
However, the analog channel inherently introduces electromagnetic interference during the transmission [12, 13, 14, 15]. While such interference enhances privacy protection, it also compromises the reliability of channel transmission, especially when it manifests as impulse interference, rendering the noise exhibiting a heavy-tailed distribution (rather than Gaussian) [16]—this has been consistently demonstrated by both theoretical[17] and empirical evidence [18]. In heavy-tailed distributions, extreme values (i.e., very large or very small values) occur with high probability, which could lead to severe signal distortion, resulting in a gradient explosion in the FL system and thereby profoundly affecting the training process of OTA FL.
@@
Numerous methods have been proposed to combat the impact of strong channel noise, ranging from channel inversion[19], phase correction[20, 21], to amplitude correction and energy estimation[22]. However, these methods only enhance the channel quality and fail to cope with the heavy tail phenomenon at the algorithmic level. Gradient norm clipping (GNC) has been proposed for resolving gradient explosion problem[23], and has also been used for heavy-tailed gradient distribution problems[24], but there is a crucial limitation: Once the data statistical structure of the gradient is altered by noise, GNC struggles to maintain its effectiveness.
@@
To enhance the robustness of OTA FL against heavy-tailed noise, we introduce a novel residual clipping technique named median anchored clipping (MAC). This method constrains the magnitude of signals received after centralization, adjusts the proportional relationships among gradients, maximizes gradient retention, and mitigates the impact of heavy-tailed interference on OTA FL. Our main contributions are summarized as follows:
@@
  • We derive the convergence rate of OTA FL gradient descent algorithm under MAC.
  • We conduct substantial experiments where the results show that our MAC algorithm effectively mitigates the impact of heavy-tailed noise in analog OTA FL.
@@ II-A Setting
We consider the FL system depicted in Fig. 1, consisting of an edge server and $N$ clients. Every client $n$ possesses a local dataset $\mathcal{D}_{n}$ that contains $m_{n}$ data samples $\{(\boldsymbol{x}_{i},y_{i})\}^{m_{n}}_{i=1}$ where $\boldsymbol{x}_{i}\in\mathbb{R}^{d},y_{i}\in\mathbb{R}$. We assume the local datasets are statistically independent from each other. The edge server orchestrates with the clients to learn a statistical model from their datasets while preserving privacy.
@@
where $f_{n}(\boldsymbol{w})$ is the local empirical risk of agent $n$. The solution of (1) is commonly known as the empirical risk minimizer, denoted by $\boldsymbol{w}^{*}=\arg\min f(\boldsymbol{w})$. And we adopt the OTA FL for model training in this paper.
@@ II-B Federated Model Training Over the Air
The general procedure of analog OTA FL is detailed in [25]. We briefly describe it in this part for completeness. Particularly, at the $k$-th round of global communication, the edge server broadcasts the global parameter $\boldsymbol{w}_{k}$ to all the clients. Then, each client $n$ calculates its local gradient $\nabla f_{n}(\boldsymbol{w}_{k})$, modulates this parameter onto the magnitude of a set of common waveforms that are orthogonal to each other, and simultaneously sends the resulting analog signals to the edge server. The edge server passes the received signal to a bank of matched filters, with each branch tuned to one of the waveform bases, and outputs the automatically aggregated (but distorted) gradient. Formally, the global gradient can be written as follows:
@@
where $\eta$ is the learning rate. Then the global parameter will be broadcasted to all clients for the next round of computations.
@@ II-C Unstable Training Performance
Normally, the above recursion is executed multiple rounds until convergence (if it occurs), upon which all participating entities have a common model close to $\boldsymbol{w}^{*}$. However, the spectrum is, by nature, a shared medium, giving rise to potentially strong co-channel interference, which typically manifests as noise during training. A notable feature of analog channel noise is its heavy-tail characteristic[18], which is manifested by frequent occurrence of impulse noise, which could lead to gradient explosion.
@@ III-A Proposed Method
Our MAC algorithm strengthens the robustness of OTA FL against strong communication noise by performing three key steps for the aggregated gradients: 1) centralization, 2) clipping, and 3) recovery. The primary concept behind the MAC algorithm is to determine a datum point for a set of gradient entries and, anchoring on this point, recalibrate the magnitudes of entries.
@@
For a vector $\boldsymbol{w}\in\mathbb{R}^{d}$, med($\boldsymbol{w}$) is the median entry of all entries of $\boldsymbol{w}$, i.e., given $\boldsymbol{w}=(w_{1},w_{2},\dots,w_{d})^{T}$,
@@
The rationale behind centralizing the global gradient at the median is that this operation minimizes the L-1 deviation of the entries (note that due to heavy-tailed noise, the L-2 deviation of the entries may be unbounded). As such, during the subsequent clipping procedure, it preserves the original information of entries that are close to the median while eliminates the extreme values introduced by the impulse noise.
2) Clipping: Based on the centralized gradient $\boldsymbol{g}_{k}$, we perform value clipping to each entry, thereby constraining the range of individual entries within a specified threshold $C$. More concretely, for a generic entry $g_{k,i}$, $i\in[d]$, we have
@@
Toward this end, we obtain a new global gradient with the detrimental effects of heavy-tailed noise effectively alleviated while retaining the useful information as much as possible. The details of this method are summarized in Algorithm 1.
III-B Convergence Analysis
Assumption 1
The objective function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $\gamma$-strongly convex, i.e., for any $\boldsymbol{w},\boldsymbol{v}\in\mathbb{R}^{d}$, it is satisfied:
$f(\boldsymbol{w})\geq f(\boldsymbol{v})+\langle\nabla f(\boldsymbol{v}),\boldsymbol{w}-\boldsymbol{v}\rangle+\frac{\gamma}{2}\|\boldsymbol{w}-\boldsymbol{v}\|^{2}.$ (8)
Assumption 2
Assumption 3
$\|\nabla f_{n}(\boldsymbol{w})\|\leq G.$ (10)
At this stage, we are ready to present the main theoretical result of this paper as the following.
@@ 17 Function $\mathbf{MAC}$($\boldsymbol{g},\mathit{C}$):
18       $\boldsymbol{m}=\text{med}(\boldsymbol{g})\times\boldsymbol{1}$
@@ 19      $\boldsymbol{g}=\boldsymbol{g}-\boldsymbol{m}$   // Centralization
21             $g_{i}=\mathrm{sgn}(g_{i})\cdot\min\left\{|g_{i}|,C\right\}$// Clipping
@@ 24      $\boldsymbol{g}=\boldsymbol{g}+\boldsymbol{m}$   // Recovery
If the learning rate is set as $\eta\leq\frac{1}{L}$, then Algorithm 1 converges as
$\mathbb{E}[\|\boldsymbol{w}_{k}\!-\!\boldsymbol{w}^{*}\|^{2}$ $]\!\leq\!\bigg{(}\!1-\eta\Big{(}\!\gamma p_{C}\!+\!\frac{\delta}{d}\left(1\!-\!p_{C}\right)\!\Big{)}\!\bigg{)}^{k}\!\|\boldsymbol{w}_{0}\!-\!\boldsymbol{w}^{*}\|^{2}$
$+\frac{\eta\left(d\left(C\!-\!2G\right)^{2}p_{C}+dC^{2}\left(1\!-\!p_{C}\right)\right)}{\gamma p_{C}+\frac{\delta}{d}\left(1\!-\!p_{C}\right)},$ (11)
where $10<\delta<$ 1 and $p_{C}$ is
$p_{C}=\mathbb{P}\{|\xi_{k,i}|\leq C-2G|\}\sim 1-\frac{\tau^{\alpha}}{C^{\alpha}}.$ (12)
@@
We omit the proof due to space limitation. ∎
@@
The result in (11) demonstrates that regardless of the heavy-tail index $\alpha$ and scale parameter $\tau$, running OTA FL in conjunction with MAC consistently achieves a linear convergence rate (where the residual error can be reduced by decreasing the learning rate). As such, the robustness of model training is substantially enhanced, making it resilient to the detrimental effects of heavy-tailed communication noise.
@@ IV-A Experiment Setup
Figure 2: Performance comparison between MAC and the baselines, (a) training ResNet-18 on the CIFAR-10 dataset, (b) ResNet-34 on the CIFAR-100 dataset, and (c) training CNN on the FEMNIST dataset.
Figure 3: Sensitivity evaluation, (a)-(c) are training ResNet-18 on the CIFAR-10 dataset while non-i.i.d sampling, respectively correspond different heavy tail index $\alpha$, scale parameter $\tau$ and clipping threshold $C$.
System Setting: We examined the effectiveness of our proposed MAC algorithm by employing the GNC as a baseline and compared their training performances under the same configurations of OTA FL systems. In our experiment, Rayleigh fading with a parameter setting of $\mu=1$ is utilized for modeling channel fading. Unless otherwise stated, we set $N=50$, $\eta=0.03$, $E=5$, $\alpha=1.5$, $\tau=0.1$, and local batch size is 10.
Dataset and Models: We assessed performance on CIFAR-10, CIFAR-100[27], and FEMNIST (which is processed in a non-i.i.d manner)[28] using ResNet-18, ResNet-34[29], and CNN architectures, respectively. As the neural networks have a multi-layer structure, the MAC algorithm is executed parallel to the gradient blocks in a layer-wise manner. We also use a Dirichlet distribution with a concentration parameter of $Dir=0.3$[30, 31] to characterize data heterogeneity.
@@ IV-B Performance Evaluation
In Fig. 2, we compare the performance of OTA FL under a variety of model training tasks, where the automatically aggregated but noisy global gradient undergoes MAC, GNC, and no post-processing (which we dub as noisy transmission), respectively. We also display the ideal scenario without any channel corruption (including fading and noise) as the performance upperbound. This figure shows that the training process is severely compromised by heavy-tailed noise, leading to fluctuations in the test accuracy curve and a marked decrease in the maximum accuracy achievable (by comparing the noisy transmission and ideal transmission). Under this circumstance, GNC encounters challenges in maintaining effective performance when the statistical structure of gradients, such as the proportional relationships among different gradient dimensions and the distribution of entries with varying magnitudes, is substantially disrupted by impulse noise. In contrast, the implementation of the MAC algorithm allows for a considerable reduction in these fluctuations, thereby significantly enhancing both test accuracy and training stability, robustfying the OTA FL system.
@@
In Fig. 3, we examine the MAC performance by varying system-related parameters. Specifically, Fig. 3(a) and (b) compare the test accuracy of OTA FL MAC under different levels of communication noise. From Fig. 3(a), we can see that by decreasing the tail index from $\alpha=1.5$ to $\alpha=1.1$, the MAC algorithm exhibits consistent stability and achieves high test accuracy. Notably, the condition $\alpha=1.1$ represents a scenario with extreme volatility in the channel noise (where $\alpha=1$ corresponds to the Cauchy distribution characterized by undefined mean and variance), whilst MAC demonstrates remarkable robustness even under such extreme noise conditions. Fig. 3(b) examines the MAC performance in the presence of ascending scale parameter $\tau$. It is noteworthy that for $\tau=0.05$, $0.1$, $0.2$, and $0.3$, the SNR (defined as ratio between useful signal power and channel noise, i.e., SNR$=\|\nabla f(\boldsymbol{w}_{k})\|^{2}/\|\boldsymbol{\xi}_{k}\|^{2}$) are $-35$ dB, $-41$ dB, $-47$ dB, and $-50$ dB, respectively. The figure evidently confirms that our MAC algorithm exhibits substantial stability under challenging channel conditions, sustaining its performance even when the SNR reaches $-50$ dB. On the other hand, Fig. 3(c) presents a comparative analysis of the performance of the MAC algorithm across various clipping thresholds. The figure reveals that variations in threshold values result in similar convergence rates, indicating that the MAC algorithm is insensitive to the clipping thresholds, demonstrating the robustness and ease of parameterization inherent to the MAC algorithm.

### v4

@@ I Introduction
In lieu of conventional centralized training methods, federated learning (FL)[1, 2, 3, 4] offers a fresh perspective on collaborative data methodologies, enabling clients to benefit from high-quality model services without compromising the security of their private data. Specifically, clients are not required to upload private local data to the server; rather, they only need to transmit model weights or gradient parameters, effectively averting the exposure of sensitive information. FL establishes a collaborative connection between clients and servers, facilitating the joint learning of a shared global model.
Despite the tremendous success of the FL paradigm, there still exist obvious problems while the executing process. The frequent transmission of model information between clients and server consumes substantial network bandwidth, while the aggregation of a large number of parameters requires extensive computing resources. These factors can result difficulty in accommodating massive access and meeting stringent latency requirements. Moreover, although FL avoids the direct aggregation of user data, model parameter information such as gradients can still potentially compromise user privacy. Inference attackers may exploit shared gradients to reverse-engineer training data, leading to potential indirect breaches of privacy [5].
@@
A viable solution to this problem is by integrating over-the-air (OTA) computation[6, 7, 8, 9] into the FL system. By leveraging the superposition property of multiple-access channels, OTA computation enables the automatic aggregation of clients’ gradients, significantly improving channel utilization while simultaneously reducing computational overhead [10]. Furthermore, as the server receives aggregated gradients rather than individual ones from clients [11], the system’s vulnerability to inference attacks is substantially mitigated.
@@
However, the analog channel inherently introduces electromagnetic interference during the transmission [12, 13, 14, 15]. While such interference enhances privacy protection, it also compromises the reliability of channel transmission, especially when it manifests as impulse interference, rendering the noise exhibiting a heavy-tailed distribution (rather than Gaussian) [16]—this has been consistently demonstrated by both theoretical[17] and empirical evidence [18]. In heavy-tailed distributions, extreme values (i.e. very large or very small values) occur with high probability, which could lead to severe signal distortion, resulting in a gradient explosion in the FL system, and thereby profoundly affecting the training process of OTA FL.
@@
Numerous methods have been proposed to combat the impact of strong channel noise, ranging from channel inversion[19], phase correction[20, 21], to amplitude correction and energy estimation[22]. However, these methods only enhance the channel quality and fail to cope with the heavy tail phenomenon at the algorithmic level. In the field of deep learning, gradient norm clipping (GNC) has been proposed to solve the gradient explosion problem[23], and has been widely discussed as a solution to the gradient heavy-tailed distribution problem[24, 25, 26], but there is a crucial limitation: Once the data statistical structure of the gradient is altered by noise, GNC struggles to maintain its effectiveness.
@@
To enhance the robustness of OTA FL against heavy-tailed noise, we introduce a novel residual clipping technique named median anchored clipping (MAC). In contrast to GNC, leveraging the stability of the median, this method constrains the magnitude of signals received after centralization, adjusts the proportional relationships among gradients, maximizes gradient retention, and mitigates the impact of heavy-tailed interference on OTA FL. Our main contributions are summarized as follows:
@@
  • We derive the convergence rate of the OTA FL gradient descent algorithm with MAC under non-convex conditions.
  • We conducted substantial experiments in which the results show that our MAC algorithm effectively mitigates the impact of heavy-tailed noise in analog OTA FL.
@@ II-A Setting
We consider the FL system depicted in Fig. 1, which consists of an edge server and $N$ clients. Every client $n$ possesses a local dataset $\mathcal{D}_{n}$ that contains $m_{n}$ data samples $\{(\boldsymbol{x}_{i},y_{i})\}^{m_{n}}_{i=1}$ where $\boldsymbol{x}_{i}\in\mathbb{R}^{d},y_{i}\in\mathbb{R}$. We assume that the local datasets are statistically independent from each other. The edge server orchestrates with the clients to learn a statistical model from their datasets while preserving privacy.
@@
where $f_{n}(\boldsymbol{w})$ is the local empirical risk of agent $n$. The solution of (1) is commonly known as the empirical risk minimizer, denoted by
$\boldsymbol{w}^{*}=\arg\min f(\boldsymbol{w}).$ (2)
In this paper, we adopt the OTA FL for model training, which employs analog transmissions for intermediate parameter uploading, harnessing the superposition property of radio signals for fast gradient aggregation.
@@ II-B Federated Model Training Over the Air
The general procedure for analog OTA FL is detailed in [27]. We briefly describe it in this part for completeness. In particular, at the $k$-th round of global communication, the edge server broadcasts the global parameter $\boldsymbol{w}_{k}$ to all the clients. Then, each client $n$ calculates its local gradient $\nabla f_{n}(\boldsymbol{w}_{k})$, modulates this parameter onto the magnitude of a set of common waveforms that are orthogonal to each other, and simultaneously sends the resulting analog signals to the edge server. The edge server passes the received signal to a bank of matched filters, with each branch tuned to one of the waveform bases, and outputs the automatically aggregated (but distorted) gradient. Formally, the global gradient can be written as follows:
@@
where $\eta$ is the learning rate. Then the global parameter will be broadcast to all clients for the next round of computations.
@@ II-C Unstable Training Performance
Normally, the above recursion is executed multiple rounds until convergence (if it occurs), upon which all participating entities have a common model close to $\boldsymbol{w}^{*}$. However, the spectrum is, by nature, a shared medium, giving rise to potentially strong co-channel interference, which typically manifests as noise during training. A notable feature of analog channel noise is its heavy-tail characteristic[18], which is manifested by frequent occurrence of impulse noise, which could lead to gradient explosion. These impulse noises cause severe distortion in the statistical structure of the gradients, including the proportional relationships between different gradient entries and the distribution of entries with varying numeric magnitudes. Traditional GNC only controls the total norm of the gradient, without addressing the underlying statistical structure. As a result, GNC cannot resolve the noise problem in analog OTA FL.
@@ III Median Anchored Clipping
Figure 2: MAC algorithm—illustrated with a multi-dimensional gradient displayed in a three-dimensional view—operates as follows: For an aggregated gradient with multiple entries, the vector-median value is subtracted from each entry, clipping is applied, and the gradient is then restored by adding the vector-median value back.
@@ III-A Proposed Method
As shown in Fig. 2, our MAC algorithm strengthens the robustness of OTA FL against strong communication noise by performing three key steps for aggregated gradients: 1) centralization, 2) clipping, and 3) recovery. The primary concept behind the MAC algorithm is to determine a datum point for a set of gradient entries and, anchoring on this point, recalibrate the magnitudes of entries.
@@
For a vector $\boldsymbol{w}\in\mathbb{R}^{d}$, med($\boldsymbol{w}$) is the median entry of all entries of $\boldsymbol{w}$, i.e., given $\boldsymbol{w}=(w_{1},w_{2},\dots,w_{d})^{\top}$,
@@
The rationale behind centralizing the global gradient at the median is that this operation minimizes the $\ell_{1}$ deviation of the entries (note that due to heavy-tailed noise, the $\ell_{2}$ deviation of the entries may be unbounded). As such, during the subsequent clipping procedure, it preserves the original information of entries that are close to the median while eliminates the extreme values introduced by the impulse noise.
2) Clipping: Based on the centralized gradient $\boldsymbol{g}_{k}$, we implement value clipping on each entry, thereby restricting the range of individual entries within a designated threshold $C$. More concretely, for a generic entry $g_{k,i}$, $i\in[d]$, we have
@@
Ultimately, by employing the MAC algorithm, we obtain a new aggregated global gradient with the detrimental effects of heavy-tailed noise effectively mitigated while retaining the useful information as much as possible. The details of this method are summarized in Algorithm 1.
@@ 17 Function $\mathbf{MAC}$($\boldsymbol{g},\mathit{C}$):
18       $\boldsymbol{m}=\texttt{med}(\boldsymbol{g})\cdot\boldsymbol{1}$
@@ 19      $\boldsymbol{g}=\boldsymbol{g}-\boldsymbol{m}$   // Centralization
21             $g_{i}=\texttt{sgn}(g_{i})\cdot\min\left\{|g_{i}|,C\right\}$// Clipping
@@ 24      $\boldsymbol{g}=\boldsymbol{g}+\boldsymbol{m}$   // Recovery
III-B Convergence Analysis
Assumption 1
The objective function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is lower bounded by a constant $f(\boldsymbol{w}^{*})$, i.e., for any $\boldsymbol{w}\in\mathbb{R}^{d}$, it is satisfied:
$f(\boldsymbol{w})\geq f(\boldsymbol{w}^{*})$ (9)
Assumption 2
Assumption 3
$\|\nabla f_{n}(\boldsymbol{w})\|\leq G.$ (11)
At this stage, we are ready to present the main theoretical result of this paper as follows.
@@ III-B Convergence Analysis
If the learning rate is set as $\eta\leq\frac{2}{L}$, and clipping threshold $C>\sqrt{2}G$, then Algorithm 1 converges as
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}$ $\left[\|\nabla f(w_{k})\|^{2}\right]\leq\frac{2(f(\boldsymbol{w}_{0})-f(\boldsymbol{w}^{*}))}{Kp_{C}(2-\eta L)\eta}$
$+\frac{1}{2}\eta^{2}dL\Big{(}p_{C}\big{(}\frac{\sqrt{2}}{2}C\!-\!G\big{)}^{2}\!+\!(1\!-\!p_{C})C^{2}\Big{)}$ (12)
where $p_{C}$ is
$p_{C}=\mathbb{P}\left\{|\xi_{k,i}|\leq\frac{\sqrt{2}}{2}C-G\right\}\sim 1-\frac{\tau^{\alpha}}{C^{\alpha}}.$ (13)
@@
Please refer to Appendix C. ∎
@@
The result in (12) demonstrates that regardless of the tail index $\alpha$ and scale parameter $\tau$, running OTA FL in conjunction with MAC consistently achieves a sublinear convergence rate (where the residual error can be reduced by decreasing the learning rate). As such, the robustness of model training is substantially enhanced, making it resilient to the detrimental effects of heavy-tailed communication noise.
@@
Figure 3: Performance comparison between MAC and the baselines, (a) and (d) training ResNet-18 on the CIFAR-10 dataset, (b) and (e) ResNet-34 on the CIFAR-100 dataset, and (c) and (f) training CNN on the FEMNIST dataset.
@@ IV-A Experiment Setup
System Setting: We evaluated the effectiveness of our proposed MAC algorithm by utilizing the GNC as a baseline and conducted a comparison of their training performance under identical configurations of OTA FL systems. Additionally, during the experiment, we performed threshold searches on both the MAC and GNC algorithms and conducted tuning to enhance performance, thereby ensuring the fairness of the comparison. In our experiment, Rayleigh fading with a parameter setting of $\mu=1$ is utilized to model channel fading. Unless otherwise stated, the following parameters will be used: number of clients $N=50$, learning rate $\eta=0.03$, local epoch $E=5$, $S\alpha S$ tail index $\alpha=1.5$, $S\alpha S$ scale parameter $\tau=0.1$, and the local batch size is 10.
Dataset and Models: We assessed performance on CIFAR-10, CIFAR-100[29], and FEMNIST (which is processed in a non-i.i.d. manner)[30] using ResNet-18, ResNet-34[31], and CNN architectures, respectively. As the neural networks have a multi-layer structure, the MAC algorithm is executed parallel to the gradient blocks in a layer-wise manner. In experiments involving non-independent and identically distributed (non-i.i.d.) data on the CIFAR-10 dataset, we use a Dirichlet distribution with a concentration parameter of $Dir=0.3$[32, 33] to characterize data heterogeneity.
Figure 4: Sensitivity evaluation, (a)–(c) are training ResNet-18 on the CIFAR-10 dataset while non-i.i.d. sampling, respectively correspond different tail index $\alpha$, scale parameter $\tau$ and clipping threshold $C$.
@@ IV-B Performance Evaluation
In Fig. 3, we present a comprehensive comparative analysis of the performance of OTA FL across diverse model training tasks. The automatically aggregated but noisy global gradient gradient, is subjected to various post-processing methodologies: MAC, GNC, and without any post-processing (termed as noisy transmission). Furthermore, the ideal scenario is illustrated, wherein channel corruption, encompassing fading and noise, is absent, serving as the upper bound of performance. The figures presented here demonstrate that the training process is severely impacted by heavy-tailed noise, leading to significant fluctuations in the test accuracy curve, a substantial reduction in the achievable maximum accuracy, and an increase in the loss function value, as shown by the comparison between noisy and ideal transmissions. In these circumstances, numerous experiments demonstrate that GNC faces difficulties in sustaining effective performance when the statistical structure of gradients, including the proportional relationships among various gradient dimensions and the distribution of entries with differing magnitudes, is significantly disrupted by impulse noise. Conversely, the application of the MAC algorithm results in a substantial mitigation of these fluctuations, thereby markedly improving test accuracy and training stability, while fortifying the robustness of the OTA FL system.
@@
In Fig. 4, we examine the performance of the MAC by altering system-related parameters. Moreover, distinct from Fig. 3, all experiments depicted in Fig. 4 are executed under non-i.i.d. conditions, thereby illustrating the universal applicability of our MAC algorithm across datasets characterized by varying degrees of data heterogeneity. Specifically, Fig. 4(a) and (b) illustrate a comparative evaluation of the test accuracy of OTA FL MAC across different levels of communication noise. As shown in Fig. 4(a), it is evident that decreasing the tail index from $\alpha=1.5$ to $\alpha=1.1$ results in the MAC algorithm maintaining consistent stability while achieving high test accuracy. Notably, the condition $\alpha=1.1$ represents a scenario with extreme volatility in the channel noise (where $\alpha=1$ corresponds to the Cauchy distribution characterized by undefined mean and variance), whilst MAC demonstrates remarkable robustness even under such extreme noise conditions. Fig. 4(b) examines the performance of the MAC algorithm under conditions of different scale parameter $\tau$. It is important to observe that for $\tau=0.05$, $0.1$, $0.2$, and $0.3$, the Signal-to-Noise Ratios (SNRs), defined as the ratio of useful signal power to channel noise (i.e., SNR$=\|\nabla f(\boldsymbol{w}_{k})\|^{2}/\|\boldsymbol{\xi}_{k}\|^{2}$), have average values of up to $-35$ dB, $-41$ dB, $-47$ dB, and $-50$ dB, respectively, after measurement. The figure clearly corroborates that our MAC algorithm exhibits considerable stability in adverse channel conditions, sustaining its performance even when the SNR reaches $-50$ dB. In alignment with our analysis, Fig. 4(a) and (b) illustrate that the MAC algorithm substantially mitigates the direct impact of tail index and scale parameters associated with heavy-tailed noise on the convergence rate of OTA FL, thereby enhancing system stability in the extreme case. On the other hand, Fig. 4(c) presents a comparative analysis of the performance of the MAC algorithm across various clipping thresholds. The figure reveals that variations in threshold values yield analogous convergence rates, indicating that the MAC algorithm exhibits insensitivity to the clipping thresholds, demonstrating the robustness and ease of parameterization characteristic to the MAC algorithm.